The diameter of a circle is 12 meters. Find the area of the circle.
Answer by checkley71(8403)

D=13 THUS RADIUS=12/2=6 METERS
AREA=PI*R^2
AREA=3.14*6^2
AREA=3.14*36
AREA=113.04 SQMETERS.